### v2

# An Efficient Gloss-Free Sign Language Translation Using Spatial Configurations and Motion Dynamics with LLMs

Abstract
Gloss-free Sign Language Translation (SLT) converts sign videos into spoken language sentences without relying on glosses, which are the written representations of signs. Recently, Large Language Models (LLMs) have shown remarkable translation performance in gloss-free methods by harnessing their powerful natural language generation capabilities. However, these methods often rely on domain-specific fine-tuning of visual encoders to achieve optimal results. By contrast, we emphasize the importance of capturing the spatial configurations and motion dynamics in sign language. With this in mind, we introduce Spatial and Motion-based Sign Language Translation (SpaMo), a novel LLM-based SLT framework. The core idea of SpaMo is simple yet effective: instead of domain-specific tuning, we use off-the-shelf visual encoders to extract spatial and motion features, which are then input into an LLM along with a language prompt. Additionally, we employ a visual-text alignment process as a lightweight warm-up step before applying SLT supervision. Our experiments demonstrate that SpaMo achieves state-of-the-art performance on three popular datasets—PHOENIX14T, CSL-Daily, and How2Sign—without visual fine-tuning  [^1] Code is available at https://github.com/eddie-euijun-hwang/SpaMo.
An Efficient Gloss-Free Sign Language Translation Using Spatial Configurations and Motion Dynamics with LLMs
Eui Jun Hwang Sukmin Cho Junmyeong Lee Jong C. Park School of Computing Korea Advanced Institute of Science and Technology {ehwa20,nelllpic,david516,jongpark}@kaist.ac.kr
1 Introduction
Sign language is a visual means of communication primarily used by Deaf communities, relying on physical movements rather than spoken words. In this paper, we tackle Sign Language Translation (SLT), focusing on converting sign videos into spoken language sentences. Early SLT methods Camgoz et al. (2020); Zhou et al. (2021a); Chen et al. (2022a, b); Zhang et al. (2023b) have primarily relied on glosses—written representations of signs using corresponding words. Glosses provide a structured form of sign language, which helps identify semantic boundaries within continuous sign sequences. This, in turn, allows the models to better comprehend the overall content of the sign videos Yin et al. (2023); Wei and Chen (2023). However, annotating glosses is a labor-intensive and time-consuming process that requires expertise in sign language. This significantly hinders the expansion of sign language datasets and limits the development of SLT methods Li et al. (2020b); Shi et al. (2022); Lin et al. (2023).
“Cold” Refer to caption “Winter”
(a) Spatial configuration
“Chair” Refer to caption “Sit”
(b) Motion dynamic
Figure 1: Visual examples of spatial configurations and motion dynamics in sign language. The images are sourced from WLASL Li et al. (2020a).
To address these limitations, there has been a shift towards gloss-free methods that rely solely on the sign videos and corresponding translations. While these methods still underperform compared to the gloss-based methods, efforts have been made to reduce the performance gap by focusing on temporal semantic structures Li et al. (2020b) and aligning visual and textual modalities Zhao et al. (2021); Yin et al. (2023); Zhou et al. (2023); Zhao et al. (2024). Recently, LLMs have demonstrated remarkable translation performance in a gloss-free setting by harnessing their powerful language generation capabilities. However, the modality gap between the continuous sign videos and discrete text poses a challenge for the LLMs in effectively understanding the sign videos. To address this, many methods fine-tune their visual encoders to be more domain-specific to sign language Wong et al. (2024); Chen et al. (2024); Rust et al. (2024); Gong et al. (2024).
That said, fine-tuning visual encoders can be resource-heavy and time-consuming, making it impractical for real-world applications, especially when considering the diversity of sign languages. This leads to an important question: Is domain-specific tuning of visual encoders necessary to achieve optimal performance in LLM-based SLT? We argue that focusing on the inherent characteristics of sign language could reduce the need for such extensive fine-tuning. First, visual encoders trained on general domains Radford et al. (2021); Oquab et al. (2023) have already proven highly effective in downstream tasks such as action recognition Huang et al. (2024); Tang et al. (2024) and video captioning Yang et al. (2023); Zhou et al. (2024). Moreover, LLMs are capable of retaining rich visual information from these general encoders in their latent space Zhang et al. (2024c). Rather than emphasizing fine-tuning, we shift our attention to the crucial roles of spatial configurations and motion dynamics in sign language.
Spatial configurations include the arranging and positioning of signs within the signing space, including hand shapes, facial expressions, and body postures. These components work together to distinguish different signs and convey their intended meanings Emmorey and Casey (1995). As shown in Figure 1(a), the signs for “cold” and “winter” both use the same handshape, with a shivering motion of the fists. The primary difference lies in the facial expressions: “cold” is typically accompanied by a tense or grimaced expression, while “winter” may feature a more neutral expression. Motion dynamics, on the other hand, involve the path, speed, and rhythm of hand movements, illustrating how movements alter the meanings of signs over time Bosworth et al. (2019). As shown in Figure 1(b), the signs for “chair” and “sit” both use the same “H” handshape and involve the interaction of both hands. However, the motion differentiates these signs: “chair” involves a repetitive tapping motion, while “sit” involves a single, smooth motion. These examples highlight the importance of the spatial configurations and motion dynamics in conveying accurate messages in sign language.
To this end, we introduce a novel gloss-free framework, Spatial and Motion-based Sign Language Translation (SpaMo). SpaMo is designed to fully exploit the spatial configurations and motion dynamics in the sign videos using off-the-shelf visual encoders, all without the need for domain-specific fine-tuning. As shown in Figure 2, the core idea is simple: We extract spatial features (spatial configurations) and motion features (motion dynamics) using two different visual encoders, and feed these into an LLM with a language prompt. Specifically, we use a pre-trained image encoder (e.g., ViT) as Spatial Encoder (SE) to individually encode each frame for its spatial features. To further refine the spatial features, we apply $S^{2}$ scaling Shi et al. (2024), which processes a sign image at multiple scales. Additionally, we use a video encoder (e.g., VideoMAE) as Motion Encoder (ME) to encode sign clips (groups of sign frames) into the motion features. To capture finer motion dynamics, we apply a sliding window approach, which results in implicit gloss-level representations Cheng et al. (2023); Hwang et al. (2024). Next, Sign Adapter (SA), comprising Multi-Layer Perceptron (MLP) layers, transfers these features to the LLM. To further bridge the modality gap, we propose Visual-Text Alignment (VT-Align), a training strategy that aligns the visual features with the LLM’s embedding space, promoting more efficient training and improved translation performance.
In all, our contributions can be summarized as:
  • We introduce SPaMo, a novel gloss-free framework based on LLMs. Our method is simple yet effective, focusing on conveying core elements of sign language to LLMs without domain-specific tuning of visual encoders.
  • Our proposed method achieves state-of-the-art performance on three popular sign language datasets: PHOENIX14T, CSL-Daily, and How2Sign.
  • We provide a novel and comprehensive analysis of how the LLM interprets the sign videos within its embedding space and translates them into corresponding text.
2 Related Work
2.1 Gloss-free Sign Language Translation
Gloss-free SLT directly converts sign videos into spoken language sentences without relying on glosses. These methods, however, often underperform compared to gloss-based methods Camgoz et al. (2020); Zhou et al. (2021b, a); Yin et al. (2021); Chen et al. (2022a, b); Zhang et al. (2023b); Jing et al. (2024). To address the performance gap, recent work has focused on several key areas: enhancing the temporal semantic structure Li et al. (2020b), improving the alignment between visual and textual modalities Zhao et al. (2021); Lin et al. (2023); Fu et al. (2023), leveraging LLMs Wong et al. (2024); Gong et al. (2024); Chen et al. (2024), and scaling efforts by utilizing larger sign language datasets Uthus et al. (2024); Rust et al. (2024). Despite these advancements, most gloss-free methods depend on fine-tuning visual encoders using the glosses Li et al. (2020b); Yin et al. (2023); Fu et al. (2023), target translations Zhou et al. (2023); Wong et al. (2024), or self-supervised learning Gong et al. (2024); Rust et al. (2024). In particular, fine-tuning with the glosses helps the visual encoders to offer more domain-specific training on continuous or isolated Sign Language Recognition (SLR) datasets, such as WLASL Li et al. (2020a) and PHOENIX14T Camgoz et al. (2018).
Consequently, we classify these methods as weakly gloss-free due to the implicit involvement of the glosses, as further elaborated in Section 4.2. On the other hand, the rest of the fine-tuning methods eliminate reliance on these annotations. However, they often require substantial resources. As a results, it can be difficult to achieve robust visual representations and improve translation performance without access to a sufficiently large dataset. To address this limitation, our approach diverges from this norm by focusing on capturing the spatial configurations and motion dynamics through off-the-shelf visual encoders. This allows us to bypass the need for resource-intensive fine-tuning.
2.2 Scaling Language Models in SLT
The scaling laws in language models Kaplan et al. (2020) have been pivotal in the rise of Large Language Models (LLMs) Touvron et al. (2023); Chiang et al. (2023); Chung et al. (2024). Leveraging their strong generation capabilities, LLMs have been applied across diverse domains: multilingual translation Zhu et al. (2023); Zhang et al. (2023a); Gao et al. (2024), pose generation Feng et al. (2024); Zhang et al. (2024b), and visual question answering Li et al. (2023); Liu et al. (2024a, b), extending their impact beyond Natural Language Processing. LLMs have also demonstrated impressive translation performance in the SLT domain. SLT methods using the LLMs focus on aligning high-dimensional visual features with inputs comprehensible to LLMs. These methods involve fine-tuning visual encoders to produce language-like tokens Gong et al. (2024), using pseudo-glosses Wong et al. (2024), or performing video-grounded text generation tasks Chen et al. (2024).
In this work, we take a different approach by focusing on spatial configurations and motion dynamics. We extract spatial and motion features and pass them to LLMs with a light warm-up process. This method is simple yet effective, demonstrating that an extensive pre-training for the visual encoders is unnecessary to achieve peak performance.
Figure 2: An overview of the SpaMo framework, which consists of three parts: (i) Sign Feature Extraction: Spatial and motion features are extracted using SE and ME, using the $S^{2}$ and sliding window approaches to capture detailed spatial configurations and motion dynamics. (ii) VT-Align: The extracted features are combined within SA to form a unified sign feature. During training, a warm-up process is employed to ensure that SA has well-initialized weights, effectively bridging the modality gap between the sign video and text. (iii) LLM: the LLM processes the sign feature along with a language-instructive prompt and is trained using LoRA.
3 Method
We first give an overview of our framework in Section 3.1. We then explain SE and ME in Sections 3.2 and 3.3, respectively. Next, we discuss SA in Section 3.4 and VT-Align in Section 3.5. Finally, we explain the training details in Section 3.6.
3.1 Framework Overview
Given a sign video $X=\{x_{i}\}_{i=1}^{T}$, where each frame $x_{i}\in\mathbb{R}^{H\times W}$ represents a frame with height $H$ and width $W$, the objective of SLT is to generate a corresponding spoken language sentence $Y=\{y_{j}\}_{j=1}^{U}$, composed of $U$ words. Previous gloss-free methods Zhou et al. (2023); Wong et al. (2024); Gong et al. (2024); Chen et al. (2024) have involved fine-tuning visual encoders using sign language data to make them more domain-specific, leading to improvements in translation accuracy. However, while this fine-tuning introduces more domain knowledge at the feature extraction level, it is unnecessary, especially with LLMs, which already maintain rich visual information from the visual encoder in their latent space Zhang et al. (2024c). Although there may be a trade-off, we argue that utilizing the spatial configurations and motion dynamics with proper alignment and training, offers a more efficient and effective solution.
As shown in Figure 2, Spatial Encoder (SE) and Motion Encoders (ME) extract two distinct features from the sign video $X$: Spatial features $Z_{s}$ capture the spatial configurations Emmorey and Casey (1995), and motion features $Z_{m}$ represent the motion dynamics Bosworth et al. (2019). These features are then integrated into a combined sign feature $Z_{sm}$ via Sign Adapter (SA). The combined feature is then fed to an LLM with a language prompt, guiding the LLM to generate the translation in the desired language. Additionally, we perform Visual-Text Alignment (VT-Align) to minimize the gap between the visual and textual modalities before and during training under SLT supervision. In the following sections, we provide a detailed explanation of SE, ME, SA, and VT-Align.
3.2 Spatial Encoder
SE extracts spatial features $Z_{s}$ from the sign video $X$. We utilize a pre-trained image encoder (e.g., ViT), which is kept frozen, and enhances its capability to capture finer spatial information by applying Scaling on Scales ($S^{2}$hi et al. (2024). $S^{2}$ is parameter-free and enables the extraction of multi-scale features without altering the original pre-trained encoder. By processing sign images at multiple resolutions, $S^{2}$ provides a more comprehensive spatial understanding, ensuring that SE captures both fine-grained and broad spatial details for accurate sign language interpretation. The resulting spatial features can be represented as $Z_{s}\in\mathbb{R}^{T\times 2d}$, where $T$ is the number of frames, and $2d$ is the increased embedding dimension, reflecting the integration of multi-scale features.
3.3 Motion Encoder
ME derives motion features from the sign video $X$. Similar to SE, we employ a pre-trained video encoder (e.g., VideoMAE), which remains frozen, to process sign clips segmented from the video. However, accurately segmenting the sign video into distinct gloss-level clips is challenging without the support of pre-trained Continuous Sign Language Recognition (CSLR) models Wei and Chen (2023). To address this limitation, we use a sliding window approach to capture implicit gloss-level representations Cheng et al. (2023); Hwang et al. (2024). Specifically, we divide the sign video into short, overlapping clips, then feed each clip into ME to extract the implicit gloss-level motion features $Z_{m}\in\mathbb{R}^{N\times d}$, where $N$ is the number of segments. The number of segments $N$ is calculated as $N=\left\lfloor\frac{T-w}{s}\right\rfloor+1$, where $T$ is the total number of frames, and $w$ and $s$ are the window size and stride, respectively. Since $Z_{m}$ is generated by processing $N$ short clips, it can also be interpreted as a sequence of $N$ clip-wise features.
Figure 3: An overview of Sign Adapter.
3.4 Sign Adapter
In the previous sections, we extracted two distinct visual features: the spatial features $Z_{s}$ and the motion features $Z_{m}$. These features differ in both their dimensions and representation, as depicted in Figure 2. To effectively integrate these features, we introduce an additional module called Sign Adaptor (SA). As shown in Figure 3, SA includes linear projection layers, a 1D CNN, and a Multi-Layer Perceptron (MLP). These components work together to integrate the spatial and motion features into a unified sign representation, denoted as $Z_{sm}$. First, the spatial and motion features are passed through linear projection layers to transform them into features with matching dimensions. Next, the 1D CNN is applied for short-term modeling of the combined features. Finally, a cross-modal MLP Liu et al. (2024a) is employed to bridge the visual and textual modalities. The resulting outputs are represented as $Z_{sm}\in\mathbb{R}^{M\times d^{\prime}}$, where $M$ represents the reduced number of frames after convolution, and $d^{\prime}$ is the dimension aligned with that of the LLM. Although SA aids in bridging the modality gap between visual and textual features during training under the SLT supervision, the gap remains. To tackle this issue, we introduce VT-Align, which will be detailed in the next section.
3.5 Visual-Text Alignment
VT-Align is a warm-up and go process designed to provide the SA module with well-initialized weights before the SLT supervision begins. This initial alignment is crucial, as it helps the model more effectively bridge the modality gap during training. To achieve this alignment, we employ a widely-used softmax-based contrastive learning approach Radford et al. (2021); Jia et al. (2021).
Specifically, given a mini-batch =$\mathcal{B}=\{(S_{1},Y_{1}),(S_{2},Y_{2}),...\}$ of sign-text pairs, the contrastive learning objective encourages the embeddings of matching pairs $(S_{i},Y_{i})$ to align closely while pushing apart the embeddings of mismatched pairs $(S_{i},Y_{j\neq i})$. Text features $Z_{t}$ are extracted from the target translation $Y_{i}$ using the LLM’s embedding layer $E_{llm}(\cdot)$. Note that only the SA module $f_{sa}(\cdot)$ is updated during this process, while $E_{llm}(\cdot)$ remains fixed to preserve the LLM’s language capabilities. The VT-Align loss function $\mathcal{L}_{vt}$ is represented as follows:
$-\frac{1}{2|\mathcal{B}|}\sum_{i=1}^{|\mathcal{B}|}\left(\overbrace{\log\frac{e^{\tau{Z_{sm}^{(i)}}\cdot z_{t}^{(i)}}}{\sum_{j=1}^{|\mathcal{B}|}e^{\tau{Z_{sm}^{(i)}\cdot Z_{t}^{(j)}}}}}^{\text{sign}\rightarrow\text{text softmax}}+\overbrace{\log\frac{e^{\tau{Z_{sm}^{(i)}}\cdot Z_{t}^{(i)}}}{\sum_{j=1}^{|\mathcal{B}|}e^{\tau{Z_{sm}^{(j)}\cdot Z_{t}^{(i)}}}}}^{\text{text}\rightarrow\text{sign softmax}}\right),$ (1)
where $Z_{sm}^{(i)}=\frac{f_{sa}(S_{i})}{\left\|f_{sa}(S_{i})\right\|_{2}}$, $Z_{t}^{(i)}=\frac{E_{llm}(T_{i})}{\left\|E_{llm}(T_{i})\right\|_{2}}$, and $\tau$ denotes a learnable temperature parameter used to scale the logits.
3.6 Training Details
Our framework is optimized in two phases: an initial warm-up phase followed by training with the SLT supervision. In the warm-up phase, we begin with a training phase the SA module using VT-Align for a designated number of steps (e.g., 4K steps). After completing the warm-up phase, we proceed to the joint training for both SA and the LLM. For fine-tuning the LLM, we utilize LoRA Hu et al. (2021), a lightweight and efficient method specifically designed for this purpose. Overall, our method is trained with a combined loss function:
$\mathcal{L}_{\textit{SpaMo}}=\mathcal{L}_{ce}+\mathcal{L}_{vt},$ (2)
where $\mathcal{L}_{ce}$ represents cross-entropy loss, and $\mathcal{L}_{vt}$ continuously manages the alignment.
PHOENIX14T CSL-Daily
Setting Methods Vis. Ft. B1 B2 B3 B4 RG B1 B2 B3 B4 RG
Gloss-based SLRT Camgoz et al. (2020) $\checkmark$ 46.61 33.73 26.19 21.32 - 37.38 24.36 16.55 11.79 36.74
BN-TIN-Transf.+SignBT Zhou et al. (2021a) 50.80 37.75 29.72 24.32 49.54 51.42 37.26 27.76 21.34 49.31
MMTLB Chen et al. (2022a) $\checkmark$ 53.97 41.75 33.84 28.39 52.65 53.31 40.41 30.87 23.92 53.25
TS-SLT Chen et al. (2022b) $\checkmark$ 54.90 42.43 34.46 28.95 53.48 55.44 42.59 32.87 25.79 55.72
SLTUNET Zhang et al. (2023b) $\checkmark$ 52.92 41.76 33.99 28.47 52.11 54.98 41.44 31.84 25.01 54.08
Weakly Gloss-free TSPNet Li et al. (2020b)$\ddagger$ $\checkmark$ 36.10 23.12 16.88 13.41 34.96 17.09 8.98 5.07 2.97 18.38
GASLT Yin et al. (2023) $\checkmark$ 39.07 26.74 21.86 15.74 39.86 19.90 9.94 5.98 4.07 20.35
ConSLT Fu et al. (2023) $\checkmark$ - - - 21.59 47.69 - - - - -
Gloss-free CSGCR Zhao et al. (2021) 36.71 25.40 18.86 15.18 38.85 - - - - -
GFSLT-VLP Zhou et al. (2023) $\checkmark$ 43.71 33.18 26.11 21.44 42.29 39.37 24.93 16.26 11.00 36.44
FLa-LLM Chen et al. (2024) $\checkmark$ 46.29 35.33 28.03 23.09 45.27 37.13 25.12 18.38 14.20 37.25
Sign2GPT Wong et al. (2024) $\checkmark$ 49.54 35.96 28.83 22.52 48.90 41.75 28.73 20.60 15.40 42.36
SignLLM Gong et al. (2024) $\checkmark$ 45.21 34.78 28.05 23.40 44.49 39.55 28.13 20.07 15.75 39.91
[…]
Table 1: Performance comparison on the PHOENIX14T and CSL-Daily datasets. “Vis. Ft.” denotes the visually fine-tuned on sign language datasets. $\ddagger$ denotes results reproduced by Yin et al. for CSL-Daily. The best results are highlighted as bold, and the second-best are underlined.
Setting Methods Modality Vis. Ft. BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE BLEURT
Weakly Gloss-free GloFE-VN Lin et al. (2023) Landmark $\checkmark$ 14.94 7.27 3.93 2.24 12.61 31.65
OpenSLT Tarrés et al. (2023) RGB $\checkmark$ 34.01 19.30 12.18 8.03 - -
Gloss-free YT-ASL-SLT Uthus et al. (2024$\dagger$ Landmark 14.96 5.11 2.26 1.22 - 29.98
SSVP-SLT Rust et al. (2024$\dagger$ RGB $\checkmark$ 30.20 16.70 10.50 7.00 25.70 39.30
FLa-LLM Chen et al. (2024) RGB $\checkmark$ 29.81 18.99 13.27 9.66 27.81 -
[…]
Table 2: Performance comparison of translation results on the How2Sign dataset. YT-ASL-SLT and SSVP-SLT (marked with $\dagger$) are reported without dataset scaling to ensure a fair comparison.
4 Experiments
4.1 Implementation Details
For SE and ME, we use CLIP ViT-L/14 Radford et al. (2021) and VideoMAE-L/16 Tong et al. (2022), respectively. To extract the spatial features, the sign images are interpolated to multiple scales, such as $224\times 224$ and $448\times 448$. For each scale, larger images are split into sub-images of regular size ($224\times 224$) and processed individually. These features from the sub-images are then pooled and concatenated with features from the original representation. For the motion features, each clip consists of 16 frames, based on the findings from Wilbur (2009), which suggests that this frame interval captures a single sign. We set the stride $s$ between consecutive clips to 8. We use FlanT5-XL Chung et al. (2024) as the LLM for PHOENIX14T and How2Sign, while mT0-XL Muennighoff et al. (2022) is used for the CSL-Daily. During the warm-up phase with VT-Align, we use 4K steps on PHOENIX14T and CSL-Daily, and 15K steps on How2Sign. Additional implementation details can be found in Appendix A.
4.2 Experimental Settings
Datasets.
We evaluated our method on three sign language datasets: PHOENIX14T Camgoz et al. (2018), CSL-Daily Zhou et al. (2021a), and How2Sign Duarte et al. (2021). PHOENIX14T is a German Sign Language dataset comprising 8,257 samples and a vocabulary of 2,887 German words. CSL-Daily is a Chinese Sign Language dataset with 20,654 samples and a 2,343 Chinese characters. How2Sign focuses on American Sign Language and includes 35,191 samples with a vocabulary of 16K English words. Detailed dataset statistics are provided in Appendix B.
Evaluation Metrics.
We report BLEU via SacreBLEU Papineni et al. (2002); Post (2018)222nrefs:1|case:mixed|eff:no|tok:13a/zh|smooth:exp|version:2.2.1 and ROUGE-L Lin and Och (2004). BLEU-n assesses translation precision by evaluating n-grams. ROUGE-L measures text similarity by calculating the F1 score based on the longest common subsequences between predicted and reference texts. We also report BLEURT Sellam et al. (2020) from the BLEURT-20 checkpoint [^3] https://huggingface.co/lucadiliello/BLEURT-20, which has been shown to correlate well with human judgments.
A Taxonomy of SLT.
In Section 2, we explored gloss-free methods, including those that incorporate gloss-supervised visual encoders. Although these approaches have traditionally been categorized as gloss-free, we argue that they should more accurately be described as weakly gloss-free due to their dependence on gloss-annotated data. This classification is detailed in Table 3.6. Specifically, methods such as TSPNet Li et al. (2020b), GASLT Yin et al. (2023), ConSLT Fu et al. (2023), GloFE-VN Lin et al. (2023), and OpenSLT Tarrés et al. (2023) rely on sign features extracted by visual encoders trained on continuous or isolated sign language recognition (SLR) datasets.
4.3 Comparison with State-of-the-Art
Results on PHOENIX14T and CSL-Daily.
We first compared our method with both gloss-based and gloss-free methods on PHOENIX14T. As shown in Table 3.6, most previous methods rely on the domain-specific fine-tuning of their visual encoders. By contrast, our method demonstrates consistent improvements across all reported metrics on PHOENIX14T without such fine-tuning. The only exception is ROUGE, where we achieved the second-best result. Specifically, the improvement on BLEU-4 is by a margin of 0.92, representing a 3.93% increase over SignLLM Gong et al. (2024). On CSL-Daily, which covers a broader range of topics than PHOENIX14T, the performance gains are even more pronounced. Our method achieved a margin increase of 4.8 in BLEU-4, reflecting a 30.41% improvement over SignLLM.
Results on How2Sign.
Next, we evaluated our method on How2Sign, which poses greater challenges than PHOENIX14T due to its broader open-domain nature, longer sign videos, and larger vocabulary. As shown in Table 3.6, our method outperformed previous methods across all reported metrics. Specifically, we achieved a 0.45 margin in BLEU-4 which represents a 4.66% improvement over Fla-LLM Chen et al. (2024). We see a performance gain in BLEURT, reaching 2.93, which is 7.46% higher than SSVP-SLT Rust et al. (2024) under the non-scaled dataset setting.
Component Metric
SE ME VT-Align B1 B2 B3 B4 RG
$\checkmark$ 46.44 33.79 26.07 21.11 42.15
$\checkmark$ 29.71 16.23 10.99 8.36 22.44
$\checkmark$ $\checkmark$ 47.59 35.05 27.34 22.26 43.92
$\checkmark$ $\checkmark$ 48.12 35.19 27.42 22.49 44.19
[…]
Table 3: Ablation study of the main component.
Models #Trainable Params #Total Params B4
w/o LLM 60.5M 60.5M 6.35
mBART-L Liu et al. (2020) 680M 680M 10.94
mT0-XL Muennighoff et al. (2022) 23.5M 3.5B 24.23
Llama-2 Touvron et al. (2023) 32.4M 7B 13.86
[…]
Table 4: Ablation study of the impact of LLM.
Method KDEs Entropy $\downarrow$
GFSLT-VLP Zhou et al. (2023) 0.32
SPaMo (Ours) 0.12
Table 5: Comparison of KDE entropy values across different embeddings. Lower entropy values indicate more confident and distinct representations.
Figure 4: Ablation study for SE and ME. $S^{2}$ represents Scaling on Scales, and $s$ denotes stride size. Note that the presented results do not include VT-Align.
Kernel Density Estimation.
To assess the quality of sign representations, following Ye et al. (2023), we employed the Kernel Density Estimation (KDE) to estimate the probability density functions of embeddings from GFSLT-VLP and our method on PHOENIX14T. Note that we reproduced GFSLT-VLP using the official code [^4] https://github.com/zhoubenjia/GFSLT-VLP. As shown in Table 5, our method produced more compact and confident representations than GFSLT-VLP. More details on the KDE process are provided in Appendix A.4.
4.4 Ablation Study
To further evaluate our method, we conducted extensive ablation experiments on PHOENIX14T, the most widely used sign language dataset. Additional results can be found in Appendix C.
Effect of Main Components.
We begin by evaluating the impact of the key components in our framework: SE, ME, and VT-Align. As shown in Table 4.3, using SE or ME individually results in lower performance, with ME performing the worst in terms of BLEU-4. However, combining SE and ME leads to an overall improvement. Notably, when VT-Align is integrated with SE, the performance rises, nearly matching Sign2GPT (22.52 vs. 22.49). The best results are achieved when all components (SE, ME, and VT-Align) are used together, yielding the highest scores across all metrics. This highlights the importance of each component in enhancing overall performance of SpaMo.
Effect of LLM.
Next, we explored the impact of different types of LLMs by replacing our model, as shown in Table 4.3. We compared five models, each with a different number of parameters: our method without pre-trained weights, mBART-L, mT0-XL, Flan-T5-XL, and Llama-2. Among these, Flan-T5-XL achieves the best performance, though it nearly ties with mT0-XL. Interestingly, despite its larger parameter count, Llama-2, which is also employed in SignLLM, does not outperform the others. This finding aligns with the observations of Zhang et al., suggesting that scaling up LMs does not always lead to better performance.
Effect of $S^{2}$ and Neighboring Gap.
Finally, we evaluated the effect of $S^{2}$ and the gap between neighboring clips on SE and ME, respectively. As shown in Figure 4, $S^{2}$ substantially improves translation performance, highlighting its effectiveness in helping SE capture more spatial details. Additionally, our analysis reveals that a stride size of 8 between neighboring clips yields the best results, suggesting that this stride size optimally aids ME in extracting the motion dynamics.
Ref: die neue woche beginnt noch wechselhaft und etwas kühler.
(the new week begins still changeable and somewhat cooler)
GFSLT-VLP: am montag wieder wechselhaft und kühler.
(on Monday again changeable and cooler)
Ours: die neue woche beginnt wechselhaft und wieder kühler.
(the new week begins changeable and again cooler)
Ref: sonst viel sonnenschein.
otherwise, a lot of sunshine.
GFSLT-VLP: im übrigen land viel sonne.
in the rest of the country, a lot of sun.
Ours: sonst viel sonnenschein.
otherwise, a lot of sunshine.
Table 6: Translation results on the test set compared to GFSLT-VLP on PHOENIX14T. Correctly translated 1-grams are highlighted in blue, while incorrect translations are marked in red.
Vis. Token: NORDWEST SONST FREUNDLICH STURDY
(NORTHWEST OTHERWISE FRIENDLY STURDY)
Gloss: NORDWEST FREUNDLICH
(NORTHWEST FRIENDLY)
Translation: richtung norden und westen ist es recht freundlich.
(Towards the north and west it is quite pleasant.)
Vis. Token: BLEIBT WIND WINTER
(REMAINS WIND WINTER)
Gloss: BLEIBEN WIND
(REMAIN WIND)
Translation: es bleibt windig.
(it remains windy.)
Vis. Token: LIEBE GUTEN ABEND SCHÖNEN
(DEAR GOOD EVENING BEAUTIFUL)
Gloss: GUT ABEND BEGRUESSEN
(GOOD EVENING GREETINGS)
Translation: guten abend liebe zuschauer.
(good evening dear viewers.)
Table 7: Comparison between visual tokens (Vis. Token) and their corresponding glosses. Words highlighted in green are exact matches, those in pink are semantic matches, and words in blue are absent in the gloss but appear in the translation.
Figure 5: The t-SNE visualization of sign features. Different colors represent features with distinct semantics, while gray points are other categories not listed.
4.5 Qualitative Analysis
Translation Results.
Table 6 presents two example translations on PHOENIX14T, comparing our method with GFSLT-VLP, the only other publicly available baseline. In the first example (top), our method provides an accurate translation, whereas GFSLT-VLP fails to capture the correct semantic meaning. In the second example (bottom), our method again produces a precise translation, while GFSLT-VLP introduces errors, resulting in incorrect information. These examples demonstrate the superior accuracy of our method in generating reliable translations. More translation examples are shown in Appendix C.4.
Visual Token Analysis.
We performed an additional analysis to explore how the LLM interprets the sign videos. For each visual feature, we identified the word with the shortest distance in the LLM’s embedding space, representing the closest match. Further details can be found in Appendix A.5. Figure 5 shows the t-SNE visualization of each sign feature mapped to the corresponding word. We observed that certain visual features align closely with specific words, which likely represents the semantic concepts that the LLM associates with these features. In other words, these words represent the LLM’s interpretation or labeling of the visual content. We refer to these mapped words as “visual tokens”. We further compared these visual tokens with the ground-truth glosses as shown in Table 7. To ensure a clearer and more accurate semantic comparison, repetitive words were removed from the visual tokens. Surprisingly, the LLM’s interpretation of the sign videos is similar to the glosses, though not perfectly aligned. This suggests that the LLM has learned to link particular video patterns with specific textual concepts, explaining why those words cluster near the visual features in the embedding space. Additionally, we found that the visual tokens capture words that are present in the translation but not in the glosses. This finding suggests that visual tokens may provide a more comprehensive representation than current glosses, potentially broadening their scope beyond what has been traditionally documented.
5 Conclusion
In this paper, we introduced SpaMo, a novel gloss-free SLT framework based on LLMs. Apart from the previous methods that rely on domain-specific fine-tuning of their visual encoders, SpaMo focuses on capturing the spatial configurations and motion dynamics, eliminating the need for resource-intensive fine-tuning. We also proposed VT-Align, a training strategy that effectively aligns and narrows the modality gap between the sign videos and target texts, enabling the transformation of the sign videos into inputs interpretable by the LLM. Our approach achieved state-of-the-art results on three popular datasets. Furthermore, we provided the first comprehensive analysis of how the LLM interprets the sign videos within its embedding space and translates them into corresponding text.
Limitations
Recently, scaling datasets Uthus et al. (2024); Rust et al. (2024) has consistently led to performance improvements, as seen with larger sign language datasets, such as Youtube-ASL Uthus et al. (2024) and BOBSL Albanie et al. (2021). While dataset scaling could also enhance our method, in this work, we focus on a constrained setting. Specifically, we use limited sign language datasets to evaluate and compare results, demonstrating the effectiveness of our method in resource-limited scenarios. Future work will involve expanding the dataset size to explore the full potential of our method and to assess its scalability and performance across more extensive and diverse datasets.
In this paper, we highlight that domain-specific fine-tuning of visual encoders is not essential for our method. However, this implies that our method relies on visual encoders pre-trained on general tasks such as action recognition and image captioning. To bridge this gap, we introduce a pre-alignment process and apply LoRA fine-tuning to the LLM. While this might appear to be a compromise, it significantly reduces the resource requirements compared to fine-tuning both the visual encoders and the LLM. Additionally, as we discussed in the previous paragraph, this limitation can be addressed as more data becomes available, allowing for improved scalability and performance over time.
Ethics Statement
Our work focuses on developing a practical framework for SLT with the goal of overcoming communication barriers faced by the Deaf and hard-of-hearing communities. Although our approach utilizes off-the-shelf visual encoders and LLMs, there is a possibility that the framework could produce unexpected or biased outputs due to the inherent limitations in the pre-trained models. However, we are optimistic that future advancements in LLMs will help mitigate these issues. We rely on open datasets such as PHOENIX14T Camgoz et al. (2018), CSL-Daily Zhou et al. (2021a), and How2Sign Duarte et al. (2021), which, while containing potentially identifiable information, present minimal concerns regarding personal privacy. Additionally, our method has been validated only on German, Chinese, and American sign languages, limiting its applicability to other sign languages. We call for future research in SLT to expand on a broader range of sign languages, promoting greater equity for the Deaf community.
Acknowledgment
This work was supported by the Institute for Information and communications Technology Promotion (IITP) grant funded by the Korea government (MSIT) (No. 2022-0-00010, Development of Korean sign language translation service technology for the deaf in medical environment).
Appendix
In this Appendix, we first provide additional implementation details in Section A. Then, Section B provides more details about the sign language dataset used in this study, including its statistics. In Section C, we present further experimental results. Finally, in Section D, we discuss the feasibility of existing Vision-Language Models (VLMs) in the SLT domain.
Appendix A More Implementation Details
A.1 Components of SpaMo.
In the SA module, we utilize two distinct linear projection layers tailored for the output feature of ME and SE. For short-term modeling, we employ a 1D CNN configured with a specific sequence of layers: $\{K5,P2,K5,P2\}$, where $K_{\sigma}$ represents a kernel size of $\sigma$, and $P_{\sigma}$ indicates a pooling layer with a kernel size of $\sigma$ Hu et al. (2023). To integrate features into the LLM’s embedding space, we leverage an MLP cross-modal connector Liu et al. (2024a), projecting the features into a 2048-dimensional space.
A.2 Prompt Template.
To focus the LLM on the SLT task, we employ a specific prompting strategy. Our prompt includes a clear instructive prompt: “Translate the given sentence into German.” Following this, we incorporate multilingual translations via a translation engine such as Google Translate [^5] https://cloud.google.com/translate, which are sampled from the training set. These translations are included to facilitate In-Context Learning (ICL) Brown et al. (2020). The prompt is structured as follows: “Translate the given sentence into German. [SRC] = [TRG].” Here, the source input (e.g., a sentence in French) serves as the foreign language example, and the corresponding response is the translation into the target language (e.g., German, as used in PHOENIX14T). An example of this prompt structure is provided in Table A.2. To ensure that the LLM does not directly access the target translations during training, we shuffle the translation samples so that they do not match the target translation. At test time, we select a translation pair from the training set to use as a reference.
Sign Video Input: [Extracted Sign Feature]
[…]
low mountain ranges.=dort sowie in den südlichen
mittelgebirgen ist bodenfrost möglich.
La helada del suelo es posible allí y en las cadenas
montañosas del sur.=dort sowie in den südlichen
mittelgebirgen ist bodenfrost möglich.
Le gel du sol est possible là-bas et dans les chaînes
de montagnes basses du sud.=dort sowie in den südlichen
mittelgebirgen ist bodenfrost möglich.
Table 8: An example of prompt used in this paper.
Visual Encoders (SE + ME) B1 B2 B3 B4 RG
DINOv2 + V-JEPA 45.67 32.94 25.27 20.35 41.32
DINOv2 + VideoMAE 47.31 34.60 26.90 21.86 42.50
CLIP + V-JEPA 47.82 34.71 26.76 21.66 43.68
[…]
Table 9: Ablation study on various combinations of visual encoders. The results are with VT-Align.
Methods B1 B2 B3 B4 RG
Ours (w/o LoRA) 46.11 32.65 24.69 19.67 42.91
Ours (w/ LoRA) 49.80 37.32 29.50 24.32 46.57
Table 10: Ablation on our method with and without LoRA.
Dataset Language #Vocab Train / Valid / Test Avg. No. Frame Gloss Domain
PHOENIX14T Camgoz et al. (2018) DGS 3K 7,096 / 519 / 642 116 $\checkmark$ Weather Forecast
CSL-Daily Zhou et al. (2021a) CSL 2K 18,401 / 1,077 / 1,176 119 $\checkmark$ Daily-life
How2Sign Duarte et al. (2021) ASL 16K 31,128 / 1,741 / 2,322 173 Instructional
Table 11: Statistics of three sign language datasets used in this work. DGS: German Sign Language; CSL: Chinese Sign Language; ASL: American Sign Language; Avg. No. Frame: average number of video frames.
A.3 Training.
For training, we use the AdamW optimizer Loshchilov and Hutter (2017), with $\beta_{1}=0.9$, $\beta_{2}=0.98$, and a weight decay of $0.01$. The learning rate schedule includes a cosine decay with a peak learning rate of 1e-4 and a linear warmup of over 10K steps, with a minimum learning rate of 5e-5. We train our model for 40 epochs, using a single NVIDIA A100 GPU, completing the entire process within 24 hours.
A.4 Evaluating Process with KDEs.
To evaluate the quality of the learned representations, we utilize Kernel Density Estimation (KDE) to estimate the probability density functions of the embeddings from GFSLT-VLP and ours. Due to different dimensionality between these methods (1,024 vs. 2,048), we run Principal Component Analysis (PCA) to reduce the number of dimensions while retaining the most significant variance components. This dimensionality reduction facilitates more efficient and stable KDE fitting. KDE can be expressed as:
$f_{\text{kde}}(\mathbf{z})=\frac{1}{nh^{d}}\sum_{i=1}^{n}K\left(\frac{\mathbf{z}-\mathbf{z_{i}}}{h}\right),$ (3)
where $\mathbf{z_{i}}$ denotes the representation points, $K$ denotes the kernel function, $h$ is the bandwidth parameter, $d$ is the dimensionality of the data, and $n$ is the number of data points.
The entropy of KDE is then calculated as:
$H=-\sum_{i=1}^{n}f_{\text{kde}}(\mathbf{z_{i}})\log f_{\text{kde}}(\mathbf{z_{i}}),$ (4)
where $H$ represents the entropy, and $f(\mathbf{z_{i}})$ are the estimated density values at the representation points.
A.5 Generating Visual Tokens
Inspired by the reverse engineering Ju et al. (2023), we first compute the Euclidean distance between the sign feature $Z_{sm}$ and the LLM’s embedding table $E_{llm}\in\mathbb{R}^{V\times d^{\prime}}$, where $V$ represents the vocabulary size. Each sign feature is then mapped to the word associated with the shortest distance in this space. This process can be expressed as $\text{dist}(Z_{sm},E_{llm})\leq\Delta$, where $\text{dist}(\cdot)$ denotes the Euclidean distance function, and $\Delta$ represents the shortest distance to $E_{llm}$ across all sign features.
Appendix B Statistics of Sign Language Datasets
Table 11 presents a comparative overview of three popular sign language datasets: PHOENIX14T, CSL-Daily, and How2Sign, each with distinct statistics and domain.
PHOENIX14T focuses on German Sign Language (DGS) within the specific domain of weather forecasting, featuring a relatively small vocabulary of 3K words and a concise average video length of 116 frames. It includes 7,096 training samples, 519 validation samples, and 642 test samples, with gloss annotations available. This dataset is tailored for domain-specific tasks, offering clear and repetitive patterns ideal for translation and recognition within weather-related contexts.
In comparison, CLS-Daily, a dataset for Chinese Sign Language (CSL), covers a broader range of topics than PHOENIX14T, spanning areas such as family life, medical care, school, banking, shopping, and social interactions. It features a vocabulary of 2K words and an average video length of 119 frames. The dataset includes 18,401 training samples, 1,077 validation samples, and 1,176 test samples, also with gloss annotations.
On the other hand, How2Sign focuses on American Sign Language (ASL) in the instructional domain. It offers a significantly larger and more diverse dataset, with a vocabulary of 16K words and an average video length of 173 frames. The dataset consists of 31,128 training samples, 1,741 validation samples, and 2,322 test samples, but lacks gloss annotations. The diversity and complexity of How2Sign make it particularly suitable for general sign language related tasks, especially those that involve understanding varied and intricate sign sequences.
Appendix C More Experiments
C.1 Effect of Visual Encoders.
We assess the effect of various combinations of visual encoders (SE & ME). Table A.2 shows four different encoders: DINOv2 Oquab et al. (2023), CLIP Radford et al. (2021), V-JEPA Bardes et al. (2024), and VideoMAE Tong et al. (2022). The results demonstrate that the combination of CLIP and VideoMAE delivers the highest performance, suggesting potential for further improvement as visual encoders continue to advance.
C.2 Effect of LoRA.
We evaluate the effect of LoRA on the LLM. As illustrated in Table 10, the LLM with LoRA demonstrates superior performance.
Methods Vis. Ft. #Trainable Params #Total Params B4
GFSLT-VLP Zhou et al. (2023) $\checkmark$ 215.6M 215.6M 21.44
Sign2GPT Wong et al. (2024) $\checkmark$ 16M 1.8B 22.52
Fla-LLM Chen et al. (2024) $\checkmark$ >705.6M* >705.6M* 23.09
SignLLM Gong et al. (2024) $\checkmark$ - >7B* 23.40
[…]
Table 12: Model parameter comparison. * denotes an estimated value due to the unavailability of public code. “Vis. Ft.” denotes to the visually fine-tuned on sign language datasets.
C.3 Parameter Comparisons
We present a comparison of various SLT methods, focusing on the presence of visual fine-tuning ("Vis. Ft."), the number of trainable and total parameters, and their performance as measured by the BLEU-4 score. As shown in Table C.2, our method, SpaMo, achieves the highest BLEU-4 score of 24.32 without the need for the visual fine-tuning. SpaMo requires 22.7M trainable parameters, which is relatively efficient compared to other methods such as GFSLT-VLP (215.6M), Sign2GPT (16M), and Fla-LLM (>705.6M). This demonstrates that SpaMo effectively balances model complexity and training efficiency to achieve superior performance without the additional step of the visual fine-tuning.
C.4 More Qualitative Results
We provide additional translation examples for PHOENIX14T, CSL-Daily, and How2Sign. As shown in Table 14, in PHOENIX14T, our method consistently delivers accurate translations, while GFSLT-VLP struggles to capture the correct semantic meaning.
In CSL-Daily, we present a comparison between glosses and visual tokens, as well as between reference translations and generated translations. As shown in Table D, most visual tokens are matched to the glosses, though they are not perfectly aligned. Notably, in the last three examples, the visual tokens include words that are missing from the glosses but appear in the reference translations.
For How2Sign, Table 16 presents translation results along with their corresponding visual tokens. Since How2Sign lacks gloss annotations, we include actual sign frames for qualitative comparison. Similar to the results on PHOENIX14T and CSL-Daily, many visual tokens in How2Sign are closely aligned with the translations. Note that although OpenSLT Tarrés et al. (2023) is the only publicly available baseline [^6] https://github.com/imatge-upc/slt_how2sign_wicv2023, we were unable to reproduce their results due to a broken link to the fine-tuned I3D features at the time of drafting.
Pairs Cosine Similarity
Glosses & Target Translations 0.8400
Visual Tokens & Target Translations 0.6781
[…]
Table 13: Cosine similarity between the visual tokens and the glosses.
Figure 6: Performance curves across epochs.
C.5 Cosine Similarity Between Visual Tokens and Glosses
We use Sentence-BERT Reimers and Gurevych (2019) to evaluate the similarity between the generated visual tokens and the ground-truth glosses from PHOENIX14T, using cosine similarity. Additionally, we assess the similarity between the visual tokens and the target translation, as well as the alignment between the glosses and the translation, which highlights varying degrees of correspondence.
As shown in Table C.4, and as we expected, the highest similarity occurs between glosses and target translations, indicating a strong semantic correspondence. By contrast, the visual tokens show lower similarity to both target translations and glosses, suggesting that they are not random. The reason for the score not being higher is likely the inclusion of unrelated words in the visual tokens compared to the actual glosses, as illustrated in Table 7.
C.6 Performance Curve Across Epochs
Figure 6 shows the performance curves of SpaMo using BLEU-4 and ROUGE over 40 training epochs on the PHOENIX14T dataset. For comparison, we note that other models such as SignLLM, Sign2GPT, and Fla-LLM are trained for 20, 100, and 75 epochs, respectively. These results highlight the progressive improvements in SpaMo’s performance as training advances, offering a detailed look at its efficiency relative to other models.
Appendix D Feasibility of Existing VLMs in SLT
Recent advancements in Vision-Language Models (VLMs) Alayrac et al. (2022); Li et al. (2023); Liu et al. (2024b); Cheng et al. (2024) have enabled LLMs to comprehend various modalities including images and videos, beyond just text. However, in the SLT domain, current VLM designs are not well-suited for processing long sequences of sign videos. As shown in Table 11, the average sign video length exceeds 116 frames, which is significantly longer than typical action recognition or video-text datasets, where sample lengths are often under 16 frames. For example, Flamingo Alayrac et al. (2022), a widely recognized vision-language model, uses 8 frames during training and only 32 frames during inference—far fewer than what is required for SLT. Similarly, VideoLlama2 Cheng et al. (2024) employs 8 frames for training. Moreover, recent LLM-based SLT methods Wong et al. (2024); Gong et al. (2024), including our method, can be classified within the VLM family but are specialized to process and capture long sign video sequences.
Ref: und nun die wettervorhersage für morgen sonntag den zwölften juli.
(and now the weather forecast for tomorrow Sunday the twelfth of July.)
GFSLT-VLP: und nun die wettervorhersage für morgen sonntag den zwölften juni.
(and now the weather forecast for tomorrow, Sunday, the twelfth of June.)
Ours: und nun die wettervorhersage für morgen sonntag den zwölften juli.
(and now the weather forecast for tomorrow Sunday the twelfth of July.)
Ref: in der nacht muss vor allem in der nordwesthälfte mit schauern und gewittern gerechnet werden die heftig ausfallen können.
(During the night, showers and thunderstorms are expected, especially in the northwest half, which could be heavy.)
GFSLT-VLP: heute nacht gibt es im norden teilweise kräftige schauer und gewitter die örtlich unwetterartig sein können.
(At night, showers and thunderstorms can be expected, especially in the northwest half, which can sometimes be strong.)
Ours: in der nacht muss vor allem in der nordwesthälfte mit schauern und gewittern gerechnet werden die mitunter kräftig sein können.
(During the night, showers and thunderstorms are expected, particularly in the northwest half, which may be heavy.)
Ref: und nun die wettervorhersage für morgen donnerstag den siebenundzwanzigsten august.
(and now the weather forecast for tomorrow, Thursday the twenty-seventh of August.)
GFSLT-VLP: und nun die wettervorhersage für morgen donnerstag den sechsundzwanzigsten august.
(and now the weather forecast for tomorrow, Thursday the twenty-sixth of August.)
Ours: und nun die wettervorhersage für morgen donnerstag den siebenundzwanzigsten august.
(and now the weather forecast for tomorrow, Thursday the twenty-seventh of August.)
Ref: am tag ist es im westen freundlich sonst sonne und dichtere wolken im wechsel hier und da fallen einzelne schauer.
(During the day it is friendly in the west, otherwise sun and denser clouds alternate, with occasional showers here and there)
GFSLT-VLP: am tag wechseln sonne und wolken einander ab im westen fallen mitunter gewittrige schauer.
(During the day sun and clouds alternate, in the west, occasional stormy showers may occur)
Ours: am tag ist es im westen freundlich mit sonne und dichteren wolken hier und da fallen schauer.
(During the day it is friendly in the west with sun and denser clouds, with occasional showers here and there)
Ref: abseits der gewittern weht der wind schwach bis mäßig an der küste frisch.
(Away from the thunderstorms, the wind blows weak to moderate, fresh at the coast.)
GFSLT-VLP: abgesehen von gewitterböen schwacher bis mäßiger an den küsten auch frischer wind
(Apart from thunderstorm gusts, weak to moderate, also fresh wind at the coasts.)
Ours: abseits der gewittern weht der wind schwach bis mäßig an den küsten auch frisch.
(Away from the thunderstorms, the wind blows weak to moderate, also fresh at the coasts.)
Ref: am sonntag im norden und an den alpen mal sonne mal wolken und ab und an schauer sonst ist es recht freundlich.
(On Sunday in the north and in the Alps sometimes sun sometimes clouds and occasional showers otherwise it is quite pleasant.)
GFSLT-VLP: am sonntag im norden an den alpen einige schauer sonst ist es recht freundlich.
(On Sunday in the north in the Alps some showers otherwise it is quite pleasant.)
Ours: am sonntag im norden und an den alpen mal sonne mal wolken und nur einzelne schauer sonst meist freundlich.
(On Sunday in the north and in the Alps sometimes sun sometimes clouds and only a few showers otherwise mostly pleasant.)
Ref: am mittwoch eine mischung aus sonne wolken und nebelfeldern im nordwesten hier und da schauer sonst ist es trocken.
(On Wednesday a mix of sun, clouds, and fog patches in the northwest; here and there showers, otherwise it is dry.)
GFSLT-VLP: am mittwoch gibt es viele wolken hier und da schauer vor allem im nordwesten bleibt es meist trocken.
(On Wednesday there will be many clouds; here and there showers, especially in the northwest, it remains mostly dry.)
Ours: am mittwoch eine mischung aus sonne wolken und nebel im nordwesten einige schauer sonst bleibt es meist trocken.
(On Wednesday a mix of sun, clouds, and fog in the northwest; some showers, otherwise it remains mostly dry.)
Ref: am tag scheint verbreitet die sonne im süden und westen bilden sich später gebietsweise quellwolken.
(During the day, the sun shines widely in the south, and later, isolated cumulus clouds form in the west.)
GFSLT-VLP: am tag scheint in der südhälfte häufig die sonne hier und da ein paar wolken.
(During the day, the sun often shines in the southern half, here and there a few clouds.)
Ours: am tag scheint verbreitet die sonne im süden und im äußersten westen tauchen hier und da ein paar quellwolken auf.
(During the day, the sun shines widely in the south, and in the far west, here and there, a few cumulus clouds appear.)
Ref: der wind weht mäßig bis frisch mit starken bis stürmischen böen im bergland teilweise schwere sturmböen im südosten mitunter nur schwacher wind.
(The wind blows moderately to freshly with strong to stormy gusts in the mountainous regions, partly severe storm gusts in the southeast, occasionally only weak wind.)
GFSLT-VLP: der wind weht mäßig bis frisch bei schauern sowie im südosten schwere sturmböen im bergland starker bis stürmböen.
(The wind blows moderately to freshly with showers, as well as severe storm gusts in the southeast, in the mountainous regions strong to stormy gusts.)
Ours: der wind weht mäßig bis frisch mit starken bis stürmischen böen auf den bergen schwere sturmböen im süden sonst schwacher wind.
(The wind blows moderately to freshly with strong to stormy gusts on the mountains, severe storm gusts in the south, otherwise weak wind.)
Ref: am montag überall wechselhaft und deutlich kühler.
(On Monday, everywhere is changeable and significantly cooler.)
GFSLT-VLP: am montag wird es wieder wechselhafter kühler.
(On Monday, it will be changeable and cooler again.)
Ours: am montag überall wechselhaft und deutlich kühler.
(On Monday, everywhere is changeable and significantly cooler.)
Ref: sonst ein wechsel aus sonne und wolken.
(Otherwise a mix of sun and clouds.)
GFSLT-VLP: ansonsten wechseln sich teilweise dichte wolken und sonne ab.
(Otherwise partially dense clouds and sun alternate.)
Ours: sonst ein wechsel aus sonne und wolken.
(Otherwise a mix of sun and clouds.)
Ref: und nun die wettervorhersage für morgen samstag den sechsundzwanzigsten januar.
And now the weather forecast for tomorrow, Saturday, the twenty-sixth of January.
GFSLT-VLP: und nun die wettervorhersage für morgen samstag den sechsundzwanzigsten dezember.
And now the weather forecast for tomorrow, Saturday, the twenty-sixth of December.
Ours: und nun die wettervorhersage für morgen samstag den sechsundzwanzigsten januar.
And now the weather forecast for tomorrow, Saturday, the twenty-sixth of January.
Ref: sonst ist es recht freundlich.
Otherwise it is quite pleasant.
GFSLT-VLP: sonst überwiegend freundlich.
Otherwise mostly pleasant.
Ours: sonst ist es recht freundlich.
Otherwise it is quite pleasant.
Table 14: Translation results on the test set compared to GFSLT-VLP on PHOENIX14T. Correctly translated 1-grams are highlighted in blue, while incorrect translations are marked in red.
{CJK*}
UTF8gbsn
Gloss: 你 小 张 什么 时间 认识
Vis. Token: 三 机场 哪里 什么 时候认识
[…]
Ours: 你 和 小张 什么 时候 认识 的?
Gloss: 椅子 他们 想 什么 时间 去 买
Vis. Token: 椅子 时候 什么 你们 他们 每天 什么 时候 旅游 女儿 考试?
[…]
Ours: 他们 想 什么 时候 去 买 椅子?
Gloss: 不是 他 去 见面 同学
Vis. Token: 不是为什么一个 看见他人看同学桌?
[…]
Ours: 不是, 他 是 去 看 同学 。
Gloss: 这 衣服 红 怎么样 这 是 新
Vis. Token: 孩子 北京衣服红色喜欢我他怎么样认识跑身体跑 个他 北京多?
[…]
Ours: 这 件 红色 的 衣服 怎么样? 是 今年  的 。
Gloss: 中午 吃 好了 要 多 吃 水果
Vis. Token: 饮料 了冷了 了老板安全 粥早 喜欢饮料吃饭 水果 吃饭
[…]
Ours: 吃 完 午饭 我 要 多 吃 点 水果 。
Gloss: 你们 吃 什么 我 请客
Vis. Token: 你们大家来什么他怕十分钟
[…]
Ours: 你们 吃饭 什么? 我 请 你们 吃饭 。
Gloss: 超市 我 要 买 椅子 你 去
Vis. Token: 超市 事 茶 怕 椅子 时候 什么 他 来 去 吗?
[…]
Ours: 我 要 去 超市 买 椅子, 你 去 吗?
Gloss: 我  驾驶  高铁  站  接  儿子
Vis. Token: 大憾怕我我们下雨大学今天
[…]
Ours: 天气 太 冷 了, 我们 去 打 篮球 吧 。
Gloss: 穿  暖  没有  我  想  不  买
Vis. Token: 衣服 饱 热情 急 没有 觉得 想 不是 不是
[…]
Ours: 这 件 衣服 不 保暖, 我 不 想 买 。
Table 15: Translation results on the CSL-Daily test set. Exact visual token matches within glosses are highlighted in green. Words highlighted in blue are not present in the glosses but appear in the translation. Correctly translated 1-grams are shown in blue.
Image: [Uncaptioned image]
Vis. Token: AGAIN SOMEONE ONE SHOW
Ref: again, one more time we’ll show it for you.
Ours: again, one more time.
Image: [Uncaptioned image]
Vis. Token: LITTLE MORE HOW
Ref: a little bit more then this maybe.
Ours: a little bit more about it.
Image: [Uncaptioned image]
Vis. Token: NOW GO TODAY TO TAKE LITTLE THREE SEVEN FOUR WEED OUT LITTLE HERE JUILLET VORSCHRIFTEN
Ref: and we’re going to take a little weed out here.
Ours: now we’re going to take a little bit of the weed out here.
Image: [Uncaptioned image]
Vis. Token: WANT TO REPEAT TWO LOOK ON YOURÄNG KISS AGE IS YOUR HORSE
Ref: you want to look at the age of your horse.
Ours: you want to take a look at the age of your horse.
Image: [Uncaptioned image]
Vis. Token: MANY PEOPLE NOT OTHER UNDERSTAND THOUGHT
Ref: many people don’t understand.
Ours: many people don’t understand that.
Image: [Uncaptioned image]
Vis. Token: I PRACTICE WHEN WITH B FOAMERS CAST WAS SO OROU CAN KNOW IF OR GROUP
Ref: i practice with the barton oaks dental group.
Ours: i practice with the barton oaks tennis team.
Ref: so, let’s keep doing the same thing with the arms.
Ours: so, let’s keep doing the same thing with the arms.
Ref: here, two, three, four, elbow and follow wherever you’re going to go, like the knee to the groin and your elbow.
Ours: here, two, three, four, follow through where you’re going to want to squeegee, woo, woo, your elbow.
Ref: my name is robert segundo and have fun.
Ours: my name is robert todd and have fun.
Ref: watch our next segment to learn more about natural beauty products.
Ours: watch our next segment and we’ll talk a little bit more about natural beauty products.
Ref: remember, be careful when doing your home remedies, and if you’re not sure, check with your local professional.
Ours: remember very carefully when doing your home remedies if you have a cell phone.
Ref: you can start to rotate your shoulders and start to get more comfortable with your feet by turning.
Ours: you can start rotating your shoulders and start getting comfortable with your five by rotating.
Ref: hi, i’m johanna krynytzky with hip expressions belly dance studio in st. petersburg, florida.
Ours: hi, i’m johanna krynytzky with hip expressions belly dance studio in st. petersburg, florida.
Ref: i’m going to show you how to do some step-touch side foot work for belly dancing.
Ours: i’m going to show you some step touch side and medium rock for belly dancing.
Table 16: Translation results on the How2Sign test set. Correctly translated 1-gram matches are highlighted in blue. Exact visual token matches within the translation are highlighted in green.